OpenStudy (anonymous):

At our school, 5 out of 7 students is involved in an after school activity. How many of the 700 students are in an after school activity at our school? A) 500 B) 550 C) 600 D) 650

OpenStudy (anonymous):

5/7 = x/700 Multiply both sides by 700, giving you 3500/7=x x=500

OpenStudy (anonymous):

a) 500